OpenStudy (bruh234):

Determine the intervals on which the function is decreasing and increasing and then find local minima and maxima. f(x) = (x-2)(x+3)

OpenStudy (naka354):

f(x) = (x-2)(x+3) f(x) increasing if f(x) > 0 (x-2)(x+3) > 0 x < -3 or x > 2 f(x) decreasing if f(x) < 0 (x-2)(x+3) < 0 x < 0 or x > -3

OpenStudy (naka354):

f(x) has a local minimum value when f'(x) = 0. f(x) = (x-2)(x+3) = x^2 + x - 6 f'(x) = 2x + 1 = 0 when x = -1/2 f(-1/2) = -25/4 so local minimum f(x) is -25/4 when x = -1/2.